# Preconditioned block encodings for quantum linear systems

### v1

@@ A Toeplitz matrix has the form:
The Toeplitz Approximate Inverse (TPAI) is inspired by the SPAI column minimisation method from Section B.2. However, it makes the further approximation that the inverse is itself a Toeplitz matrix. This means that only one $s\times s$ system needs to be inverted instead of $N$. More importantly from a quantum perspective, the TPAI can be encoded with circuit depth of $\mathcal{O}(d)$ where $d$ is the number of diagonals, rather than $\mathcal{O}(dN)$ for SPAI.
@@
Where $D$ is the diagonal scaling matrix described in Appendix A. For the CFD matrices considered here, this creates ones along the main diagonal and tends to equalise the entries along the off-diagonals. $S_{diag}$ is a function that takes the average value along each diagonal and uses the result for the value along the corresponding Toeplitz diagonal. This approach is only appropriate to matrices that have a predominantly diagonal sparsity pattern.
@@ Appendix D Circulant Approximate Inverse
The preconditioned system is $C^{-1}Ax=C^{-1}b$. Hence, an efficient way to compute $C^{-1}$ is needed. From Equation 34 and Equation 35:
@@ E.1 Digitising the matrix entries
To avoid rounding non-zero values to zero, the digitisation algorithm creates bins based on percentages of the mean value within each bin. The principle is illustrated in Figure 13.
Figure 13: Bin selection for digitising the matrix entries: blue = entries in ascending order, yellow = $+5\%$, grey = $-5\%$. Box indicates the extent of the bin. - all blue values in the bin are reset to the mid-point value.
The algorithm initially applies a convolution operator to the ordered state amplitudes. This creates a number of potential bins, all of which meet the criteria that the values in each bin differ by no more than a given percentage, e.g. $10\%$ as shown in Figure 13. At this stage, there are a number of overlapping bins. A marching algorithm then eliminates any overlaps by retaining bins with the largest number of entries. As with most marching algorithms, it is a greedy algorithm and is not guaranteed to produce the minimum number of bins and/or the maximum number of repeats.
@@ E.2 Collapsing equal angled rotation gates
The collapsing of equal-valued rotations is shown in Figure 14.
@@ E.3 Performance
The performance of filtering and collapsing algorithm is illustrated in Figure 15. In this example, the bin contains 56 equal-valued multiplexed rotations, all of which are a Hamming distance of 1 away from at least one other rotation. The collapsing algorithm reduces this to 34 rotations. Examination of Figure 15(a) shows that there are 16 rotations that have not been modified and none are within a Hamming distance of 1 of the other. Note that while the matrix entries have been reordered to be monotonic, this does not mean that controls correspond to a contiguous set of integers.
For larger bins, the relative performance is seen to improve. For a bin of 145 equal-valued rotations, the collapsing algorithm reduces this to 45. As the size of the matrix increases, so does the size of the bins resulting in the greater percentage reductions shown in Figure 9.
@@
Figure 16 shows the circuit with the multiplexed rotations trimmed according to the tree in Figure 15(b). The labels on the circuit gates correspcond to paths through the tree.
Figure 16: Section of matrix loading circuit after trimming. All $R_{y}$ gates have the same rotation angle. The gate indices correspond to the path labels in Figure 15(b)

### v3

@@ A Toeplitz matrix has the form:
The Toeplitz Approximate Inverse (TPAI) is inspired by the SPAI column minimisation method from Section˜B.2. However, it makes the further approximation that the inverse is itself a Toeplitz matrix. This means that only one $s\times s$ system needs to be inverted instead of $N$. More importantly from a quantum perspective, the TPAI can be encoded with circuit depth of $\mathcal{O}(d)$ where $d$ is the number of diagonals, rather than $\mathcal{O}(dN)$ for SPAI.
@@
Where $D$ is the diagonal scaling matrix described in Appendix˜A. For the CFD matrices considered here, this creates ones along the main diagonal and tends to equalise the entries along the off-diagonals. $S_{diag}$ is a function that takes the average value along each diagonal and uses the result for the value along the corresponding Toeplitz diagonal. This approach is only appropriate to matrices that have a predominantly diagonal sparsity pattern.
@@ Appendix D Circulant Approximate Inverse
The preconditioned system is $C^{-1}Ax=C^{-1}b$. Hence, an efficient way to compute $C^{-1}$ is needed. From Equation˜34 and Equation˜35:
@@ E.1 Digitising the matrix entries
To avoid rounding non-zero values to zero, the digitisation algorithm creates bins based on percentages of the mean value within each bin. The principle is illustrated in Figure˜14.
Figure 14: Bin selection for digitising the matrix entries: blue = entries in ascending order, yellow = $5\%$, grey = $5\%$. Box indicates the extent of the bin. - all blue values in the bin are reset to the mid-point value.
The algorithm initially applies a convolution operator to the ordered state amplitudes. This creates a number of potential bins, all of which meet the criteria that the values in each bin differ by no more than a given percentage, e.g. $10\%$ as shown in Figure˜14. At this stage, there are a number of overlapping bins. A marching algorithm then eliminates any overlaps by retaining bins with the largest number of entries. As with most marching algorithms, it is a greedy algorithm and is not guaranteed to produce the minimum number of bins and/or the maximum number of repeats.
@@ E.2 Collapsing equal angled rotation gates
The collapsing of equal-valued rotations is shown in Figure˜15.
@@ E.3 Performance
The performance of filtering and collapsing algorithm is illustrated in Figure˜16. In this example, the bin contains 56 equal-valued multiplexed rotations, all of which are a Hamming distance of 1 away from at least one other rotation. The collapsing algorithm reduces this to 34 rotations. Examination of Figure˜16(a) shows that there are 16 rotations that have not been modified and none are within a Hamming distance of 1 of the other. Note that while the matrix entries have been reordered to be monotonic, this does not mean that controls correspond to a contiguous set of integers.
For larger bins, the relative performance is seen to improve. For a bin of 145 equal-valued rotations, the collapsing algorithm reduces this to 45. As the size of the matrix increases, so does the size of the bins resulting in the greater percentage reductions shown in Figure˜9.
@@
Figure˜17 shows the circuit with the multiplexed rotations trimmed according to the tree in Figure˜16(b). The labels on the circuit gates correspcond to paths through the tree.
Figure 17: Section of matrix loading circuit after trimming. All $R_{y}$ gates have the same rotation angle. The gate indices correspond to the path labels in Figure˜16(b)